Name: 1. A company made the following merchandise purchases and sales during the current month: July 1 July 5 July 14 July 30 purchased 400 purchased 270 purchased 300 purchased 250 units at units at units at units at $15 each $20 each $25 each $30 each There was no beginning inventory. The Company had 700 units remaining at July 31. What is the Cost of Goods Sold under FIFO? What is the Cost of Goods Sold under LIFO? What is the Cost of Goods Sold under Weighted Average? United States)
Answer #1

Units available for sale = 400 + 270 + 300 + 250 = 1,220

Units sold = Units available for sale - Ending inventory

= 1,220 - 700

= 520

Under the First in first out (FIFO) method of inventory valuation, Cost of goods sold consists of the units from beginning inventory and earliest purchases. Ending inventory consists of the units from recent purchases.

Cost of goods sold of 520 units consists of 400 units from July 1 purchases and 120 units from July 5 purchases.

Cost of goods sold = (400*$15) + (120*$20)

= $6,000 + $2,400

= $8,400

Under the Last in first out (LIFO) method of inventory valuation, Cost of goods sold consists of the units from recent purchases. Ending inventory consists of the units from beginning inventory and earliest purchases.

Cost of goods sold of 520 units consists of 250 units from July 30 purchases and 270 units from July 14 purchases.

Cost of goods sold = (250*$30) + (270*$25)

= $7,500 + $6,750

= $14,250

Cost of units available for sale = (400*$15) + (270*$20) + (300*$25) + (250*$30)

= $6,000 + $5,400 + $7,500 + $7,500

= $26,400

Cost per unit under the Weighted average method = Cost of units available for sale / Number of units available for sale

= $26,400 / 1,220

= $21.64

Cost of goods sold = 520 units * $21.64

= $11,253